A father who is 42 years old has a son and a daughter. The daughter is three times as old as the son. In 10 years the sum of all their ages will be 100 years. How old are the two siblings at present?
Answer by Cromlix(4381)

Hi there,
Make son's age = 'x'
Daughters age = 3x
Father's age 42 (given)
In 10 years all their ages total 100.
Father (42 + 10)
Father (52)
Daughter (3x + 10)
Son (x + 10)
52 + 3x + 10 + x + 10 = 100
Collect like terms:
3x + x = 100 - 52 - 10 - 10
4x = 28
x = 7
Son is 7 years old
Daughter is 21 years old.
Hope this helps :-)